# Joint Modelling of Astrophysical Systematics for Cosmology with LSST

### v2

@@ 3 Forecasting set-up
Figure 3: The redshift evolution of the fiducial IA amplitude $A(L,z)$. The Y1 case from this work (‘JMAS Y1’) is represented by a green solid line, contrasting with the JMAS Y10 scenario’s blue solid line. Dashed lines illustrate the SRD cases for comparison. Notably, the JMAS Y1 IA amplitude exceeds that of Y10, primarily because the latter captures a larger fraction of fainter galaxies, which typically exhibit weaker intrinsic alignments. Both JMAS and SRD IA amplitudes display similar distribution peaks, with minor discrepancies due to differences in redshift distribution and luminosity function parameters. The IA amplitude calculation factors in the fraction of red galaxies, leading to divergence between the JMAS and SRD scenarios. An inset visualizes this: Y1 forecast (green) and Y10 (blue) for JMAS are in solid lines, while SRD results are dashed. The red galaxy fraction in the JMAS case peaks earlier, and overall amplitudes are lower, influenced by the adjusted luminosity function parameters affecting galaxy density $P$, rate of change of the pivot luminosity $Q$, and the faint end slope $\alpha$.
@@
Data vector. Our data vector consists of auto- and cross-correlation angular power spectra $C_{\ell}$ of cosmic shear. Considering there are 5 bins (Fig. 2) in both Y1 and Y10, there are 15 angular power spectra in total. For the multipole range, we adhere to the guidelines of Mandelbaum et al. 2018 and utilize a set of 20 $\ell$ bins, defined globally and spaced equally on a logarithmic scale, within the range $00020\leq\ell\leq 15,$ 000. Scale cuts were applied in the same manner as it was done in the Mandelbaum et al. 2018, setting all values for $\ell>3000$ to zero. We do not model baryons in our analyses. As with the redshift distribution, we employ a validation step and compare the jointly modeled data vectors to those of the standard analysis (fiducial data vectors from The LSST Dark Energy Science Collaboration et al. 2021). In this process, we meticulously examine the ratio of the power spectra ${C^{ij}_{\mathrm{JMAS}}(\ell)}$ and ${C^{ij}_{\mathrm{SRD}}(\ell)}$ for each correlation, performing this analysis separately for both forecasting years 1 and 10. Specifically, we calculate the relative difference between the JMAS and SRD data vectors for each correlation (for both forecasting scenarios) to assess the relative deviations. Our findings are reassuring; the deviations do not exceed $15\%$ in any case, indicating a high degree of consistency between the jointly modeled data vector and that from the standard analysis. This outcome, shown in Fig. 12, not only validates our modeling approach but also underscores the robustness of our forecasts, as the minimal deviations suggest that the JMAS model aligns closely with the established SRD benchmarks.
@@
Covariance matrix. To ensure consistency, we have utilized the covariance matrices from the SRD study, as detailed in (Mandelbaum et al., 2018). These matrices, along with other SRD resources, are publicly available and can be found in (The LSST Dark Energy Science Collaboration et al., 2021). For an in-depth understanding of the construction and characteristics of these covariance matrices, readers are referred to the comprehensive descriptions in (Mandelbaum et al., 2018) and the original work by (Krause & Eifler, 2017).
@@
Fisher forecasting infrastructure. To perform Fisher analysis, we have developed a specialized code, FISK (Fisher Inference for Systematics in Kosmology). This code is designed as an object-oriented pipeline, incorporating a range of structural elements that align with the key components discussed in earlier sections of this document. At its core, FISK consists of various classes corresponding to each model type, including NZ (Redshift Distribution), IA (Intrinsic Alignment), LF (luminosity function), and a cosmological model. These classes contribute to the construction of the primary observable, the cosmic shear data vector using pyCCL. Based on user input, the pipeline can execute a standard analysis (SRD) or follow the path of joint modeling (JMAS). Additionally, as we will demonstrate, there is a third option (SRD+LF) that maintains the SRD framework while incorporating derivatives over LF parameters. The calculation of derivatives of the data vector with respect to model parameters is handled by a dedicated ’Derivative’ class. Lastly, the calculated derivatives are integrated with the inverse of the covariance matrix, resulting in the construction of the Fisher matrix. The FISK algorithm is illustrated in Fig. 4. The dashed line in Fig. 4 shows the route of the JMAS framework while the SRD standard analysis is depicted with dotted line. Solid line represents the mode common for both forecasting modes (SRD and JMAS). The user can also specify the forecasting year (1 or 10) or, if desired, change the value of the fiducial parameters for all the models of the study.
@@
As noted in Sec. 2.6, we encountered numerical instabilities when calculating derivatives using the five-point stencil or methods available in the numdifftools (Brodtkorb & D’Errico, 2015) Python library. To address the limitations of traditional finite difference techniques, we implemented a stem method, as detailed in (Camera et al., 2016). Following their approach, we calculate the data vector $C^{ij}(\ell)$ at 15 points surrounding each model fiducial parameter $\theta_{\alpha}$: $\delta\theta_{\alpha}=0,\pm 0.625\%,\pm 1.25\%,\pm 1.875\%,\pm 2.5\%,\pm 3.75\%,\pm 5\%,\pm 10\%$. Considering the anticipation of linear results, we strive to maintain the fit’s deviation from the actual values within our targeted precision of 1%. Should this level of precision not be met, edge data points are excluded, and the fitting process iterated until the desired precision is attained. Upon fulfilling the precision criteria, we define the derivative as the slope of the linear fit. This method has proven highly effective and stable against variations in redshift range resolution or other factors influencing the angular power spectra. Our thorough testing of this derivative calculation approach (see Figs. 15 and 16) demonstrates its robustness and superiority over the finite difference method for the analysis case we present here. Additionally, we have conducted extensive validations of our Fisher matrices by comparing them with the Fisher matrices released in the LSST DESC Science Requirements Document (The LSST Dark Energy Science Collaboration et al., 2021).
In addition to its core functionality, FISK is thoughtfully designed to allow for the generation of intermediate results at any stage of the pipeline. This feature is particularly beneficial for validation purposes, as it enables thorough examination and verification of each step in the process. Furthermore, the code is meticulously crafted to be easily extendable, modifiable, and upgradable. This flexible architecture ensures that FISK can readily adapt to evolving investigation, accommodate new models or methodologies, and integrate advancements in future analyses (for example, including the modelling of the galaxy bias). Such a design approach not only enhances the utility and longevity of FISK but also fosters a dynamic environment for continuous improvement and innovation in Fisher analysis.
@@ 4.1 Understanding the impact of the luminosity function parameters on the redshift distribution
We start by examining how changes in the parameters of the luminosity function affect the redshift distribution using the convolution method described in Section 2.2. Understanding these effects will help us interpret later results and create realistic LSST-like redshift distributions as outlined in Section 3.
In our comprehensive analysis, we employed the convolution methodology within the LSST survey framework. Our objective was to elucidate the intricate influences of individual Schechter luminosity function parameters on the redshift distribution, a critical aspect for understanding results of our Fisher forecast further on. Each parameter was varied in isolation, with others held constant at their central values, enabling a focused investigation into the unique effects of each factor. We emphasize that in this framework, the survey volume is kept constant at all times, as explained in Sec. 3 on redshift distribution modelling. The findings are shown in Fig. 5.
@@
    The normalization factor, $\phi_{0}^{*}$, which denotes the number density of galaxies at a benchmark luminosity and redshift, predominantly modulates the overall amplitude of the redshift distribution. This parameter acts as a scaling factor, and its variation is indicative of changes in the overall galaxy population density over cosmic time. The parametric sweep over the normalisation parameter is indicated in green solid lines in the Fig. 5. An increase in $\phi_{0}^{*}$ suggests a denser universe with a higher overall number of galaxies, amplifying the amplitude of the distribution.
@@
    In the context of the luminosity function’s influence on the redshift distribution of galaxies, the parameter $P$, which encapsulates the concept of density evolution within the galactic ensemble, exerts a multifaceted impact. Specifically, $P$ plays a pivotal role in modulating both the amplitude and the peak positioning of the redshift distribution. An increase in the $P$ value indicates a heightened density of galaxies at higher redshifts, as illustrated by the hot pink solid lines in Figure 5. This rise in density not only boosts the total count of galaxies within the observed distribution but also leads to a notable shift in the distribution’s peak towards higher redshifts. Such a shift, facilitated by a higher $P$ value, describes a more compact arrangement of galaxies, leading to lower lacunarity within our fixed volume element. Consequently, this decreased lacunarity at higher redshifts denotes a denser and more uniform presence or detectability of galaxies, marking a significant evolution in the cosmic structure and emphasizing the intricate relationship between galaxy density and the distribution’s overall uniformity. This nuanced interplay highlights the criticality of considering density evolution parameter in the modeling and interpretation of distribution of galaxies across various redshifts. However, the impact of a positive $P$ value, particularly in the construction of the redshift distribution, merits careful attention. With a positive $P$ the normalization factor $\phi^{*}$ increases exponentially with redshift, which can lead to a mathematical ’blow-up’ of the redshift distribution at higher redshifts. This phenomenon is critical to monitor as it may result in a redshift distribution that significantly deviates from empirical observations and established cosmological theories. Therefore, understanding and accounting for the influence of a positive $P$ is essential in accurately modeling and interpreting the distribution of galaxies across various redshifts.
@@
    In the analysis exploring the interplay between the luminosity function and the redshift distribution of galaxies, a key observation centers around the "knee" of the luminosity function, characterized by the parameter $M_{0}^{*}$. We have systematically increased $M_{0}^{*}$ from -22 to -18 in increments of 1, as shown in blue solid lines in Fig. 5. Notably, as the characteristic magnitude increased (indicating a shift towards dimmer characteristic luminosities), there was a marked decrease in the amplitude of the galaxy redshift distribution. Concurrently, a shift of the peak of the redshift distribution towards lower redshifts with larger $M_{0}^{*}$ values is present. This suggests that galaxies around and below the "knee" (dimmer galaxies) are more prevalent or detectable at lower redshifts, significantly influencing the shape and peak of the redshift distribution. These observations can help us in understanding how the properties of the "knee" of the luminosity function, especially the characteristic magnitude, affect the overall distribution and observation of galaxies across different redshifts.
@@
    When modeling the redshift distribution of galaxies based on the luminosity function, an increase in the (fiducially negative) $\alpha$ value (indicating a flatter faint-end slope and hence a greater number of low-luminosity galaxies) is observed to cause a decrease in the amplitude of the redshift distribution (distributions in purple in Fig. 5). This agrees with what we know to be true for an apparent-magnitude-limited observing set-up: that as the population of intrinsically dimmer galaxies becomes more prevalent (with more negative $\alpha$ values), the overall observed number of galaxies is lower. Additionally, there is a noted shift in the peak of the redshift distribution towards higher redshifts as $\alpha$ increases towards zero. This shift could be attributed to an increased relative representation of higher-luminosity galaxies (still in the faint-end i.e. below pivot luminosity), recalling that at higher redshifts we are preferentially going to observer intrinsically brighter galaxies.
In summary, these parameters collectively sketch a dynamic portrait of the galaxy population, each contributing to the shape of the redshift distribution of galaxies over cosmic time. Furthermore, it is evident from the mathematical expressions (Eq. 4) that there exists a degree of degeneracy among these parameters. Such degeneracies are critical in understanding the limitations and constraints of our models and interpretations. The exact nature and directions of these degeneracies will be thoroughly explored and delineated in the forthcoming discussion on Fisher analysis results. The Fisher analysis will provide a quantitative framework to dissect and understand these intricate relationships, offering a more nuanced and precise understanding of how these parameters interact and influence our results.
@@
Figure 6: Impact of Luminosity Function Parameters on Cosmic Shear Angular Power Spectra. This figure illustrates the relative difference in cosmic shear angular power spectra $C_{\ell}$, as a function of various Schechter luminosity function parameters. LF parameters are varied over five values, with the central value serving as the reference for relative comparison. In each plot, only the first bin correlations are displayed (bin 1- bin 1 correlation). The normalization factor $\phi_{0}^{*}$ is shown in green solid lines, highlighting that an increase leads to a positive relative difference ratio, indicative of a denser galaxy field enhancing the cosmic shear signal. The density evolution parameter $P$ is depicted in hot pink solid lines, where an increase results in a higher density of galaxies at higher redshifts, thus augmenting the shear signal. The characteristic magnitude $M_{0}^{*}$ is in blue, and the rate of change $Q$ in orange-yellow, both showing a similar positive trend upon increase. Conversely, an increase in the faint-end slope $\alpha$ (in purple) is observed to also contribute to the shear signal, evident from the deviation pattern across different scales. The adjustment of each luminosity function parameter, though executed in equal incremental steps, does not uniformly translate to proportional changes in the amplitude of $C_{\ell}$. In some instances it is clear that as we deviate further from the central value, either by increasing or decreasing a parameter, the impact on the amplitude changes disproportionately. This non-linear response highlights the complex, non-proportional relationship between the variations in luminosity function parameters and their consequent effects on the cosmic shear strength. These variations collectively underscore the complex interplay between LF parameters and cosmic shear, offering insights into the galaxy distribution’s influence on gravitational lensing.
@@
    Increasing the normalisation factor $\phi_{0}^{*}$, resulted in the relative difference ratio consistently being below zero (left top panel in Fig. 6). Physically, a higher $\phi_{0}^{*}$ implies an increased number density of galaxies with respect to luminosity We attribute this effect to the impact of the LF parameter on the IA’s contribution to the spectrum. As a normalisation factor, the effect of $\phi_{0}^{*}$ will ultimately be normalised out of the lensing contribution. However, due to the way it is incorporated into the IA amplitude in our model, this normalisation does not occur. So, an increase in $\phi_{0}^{*}$ increases the IA contribution and thus causes a relative decrease in the overall spectrum.
@@
    Similarly, an increase in the density parameter $P$, led to an augmented relative difference ratio, as shown in the top middle panel in Fig. 6. This result can be interpreted as being primarily due to the lensing contribution to the spectrum, as increasing $P$ shifts the mean redshift of the tomographic bin up and therefore increases the amplitude of the lensing signal.
@@
    An increase in (fiducially negative) $\alpha$, the faint-end slope, also increases the shear signal, as can be seen in the bottom right panel in Fig. 6. This indicates that a shallower faint-end slope, corresponding to a relative higher number of high-luminosity galaxies, ultimately contributes to the overall cosmic shear signal (and increases its amplitude).

### v3

@@ 3 Forecasting set-up
Figure 3: The redshift evolution of the fiducial IA amplitude $A(L,z)$. The IA amplitude is calculated using the luminosity function with the fiducial parameter values provided in Table 2. The Y1 case from this work (‘JMAS Y1’) is represented by a green solid line, contrasting with the JMAS Y10 scenario’s blue solid line. Dashed lines illustrate the SRD cases for comparison. Notably, the JMAS Y1 IA amplitude exceeds that of Y10, primarily because the latter captures a larger fraction of fainter galaxies, which typically exhibit weaker intrinsic alignments. Both JMAS and SRD IA amplitudes display similar distribution peaks, with minor discrepancies due to differences in redshift distribution and luminosity function parameters. The IA amplitude calculation factors in the fraction of red galaxies, leading to divergence between the JMAS and SRD scenarios. An inset visualizes this: Y1 forecast (green) and Y10 (blue) for JMAS are in solid lines, while SRD results are dashed. The red galaxy fraction in the JMAS case peaks earlier, and overall amplitudes are lower, influenced by the adjusted luminosity function parameters affecting galaxy density $P$, rate of change of the pivot luminosity $Q$, and the faint end slope $\alpha$.
@@
Data vector. Our data vector consists of auto- and cross-correlation angular power spectra $C_{\ell}$ of cosmic shear. Considering there are 5 bins (Fig. 2) in both Y1 and Y10, there are 15 angular power spectra in total. For the multipole range, we follow Mandelbaum et al. (2018) in considering only $\ell\leq 3000$. Our particular choice of $\ell$ bins was constructed to match that of Mandelbaum et al. (2018), to enable us to take advantage of data products from that work, particularly covariance matrices. Specifically: 20 $\ell$ bins are defined to be logarithmically spaced within the range $00020\leq\ell\leq 15,$ 000, prior to then performing the above cuts to restrict to $\ell\leq 3000$. We do not model baryons in our analyses. As with the redshift distribution, we employ a validation step and compare the jointly modelled data vectors to those of the standard analysis (fiducial data vectors from The LSST Dark Energy Science Collaboration et al. 2021). In this process, we examine the ratio of the power spectra ${C^{ij}_{\mathrm{JMAS}}(\ell)}$ and ${C^{ij}_{\mathrm{SRD}}(\ell)}$ for each correlation, performing this analysis separately for both forecasting years 1 and 10. Specifically, we calculate the relative difference between the JMAS and SRD data vectors for each correlation (for both forecasting scenarios) to assess the relative deviations. Our findings are reassuring; the deviations do not exceed $15\%$ in any case, indicating that it is possible to achieve a high degree of consistency between data vectors modelled in our framework and those calculated within a more standard modelling set-up.
@@
Covariance matrix. To ensure consistency, we have utilized the covariance matrices from the SRD study, as detailed in Mandelbaum et al. (2018). These matrices, along with other SRD resources, are publicly available and can be found in The LSST Dark Energy Science Collaboration et al. (2021). For an in-depth understanding of the construction and characteristics of these covariance matrices, readers are referred to the comprehensive descriptions in Mandelbaum et al. (2018) and the original work by Krause & Eifler (2017).
@@
Fisher forecasting infrastructure. To perform Fisher analysis, we have developed a specialized code, FISK (Fisher Inference for Systematics in Kosmology). This code is designed as an object-oriented pipeline, incorporating a range of structural elements that align with the key components discussed in earlier sections of this document. At its core, FISK consists of various classes corresponding to each model type, including NZ (Redshift Distribution), IA (Intrinsic Alignment), LF (luminosity function), and a cosmological model. These classes contribute to the construction of the primary observable, the cosmic shear data vector using pyCCL. Based on user input, the pipeline can execute a standard analysis (SRD) or follow the path of joint modelling (JMAS). Additionally, as we will demonstrate, there is a third option (SRD+LF) that maintains the SRD framework while incorporating derivatives over LF parameters. The calculation of derivatives of the data vector with respect to model parameters is handled by a dedicated ‘Derivative’ class. Lastly, the calculated derivatives are integrated with the inverse of the covariance matrix, resulting in the construction of the Fisher matrix. The FISK algorithm is illustrated in Fig. 4. The dashed line in Fig. 4 shows the route of the JMAS framework while the SRD standard analysis is depicted with a dotted line. The solid line represents the mode common for both forecasting modes (SRD and JMAS). The user can also specify the forecasting year (1 or 10) or, if desired, change the value of the fiducial parameters for all the models of the study.
@@
As noted in Sec. 2.6, we encountered numerical instabilities when calculating derivatives using the five-point stencil or methods available in the numdifftools (Brodtkorb & D’Errico, 2015) Python library. To address the limitations of traditional finite difference techniques, we implemented a stem method, as detailed in Camera et al. (2016). Following their approach, we calculate the data vector $C^{ij}(\ell)$ at 15 points surrounding each model fiducial parameter $\theta_{\alpha}$: $\delta\theta_{\alpha}=0,\pm 0.625\%,\pm 1.25\%,\pm 1.875\%,\pm 2.5\%,\pm 3.75\%,\pm 5\%,\pm 10\%$. For parameters with a fiducial value of zero (e.g., $w_{a}$), we use the percentage changes directly, without scaling by $\theta_{\alpha}$, ensuring the method remains robust even in such cases. Considering the anticipation of linear results, we strive to maintain the fit’s deviation from the actual values within our targeted precision of 1%. Should this level of precision not be met, edge data points are excluded, and the fitting process iterated until the desired precision is attained. Upon fulfilling the precision criteria, we define the derivative as the slope of the linear fit. This method has proven highly effective and stable against variations in redshift range resolution or other factors influencing the angular power spectra. Our thorough testing of this derivative calculation approach (see Figs. 16 and 17) demonstrates its robustness and superiority over the finite difference method for the analysis case we present here. Additionally, we have conducted extensive validations of our Fisher matrices by comparing them with the Fisher matrices released in the LSST DESC Science Requirements Document (The LSST Dark Energy Science Collaboration et al., 2021).
In addition to its core functionality, FISK is thoughtfully designed to allow for the generation of intermediate results at any stage of the pipeline. This feature is particularly beneficial for validation purposes, as it enables thorough examination and verification of each step in the process. Furthermore, the code is designed to be easily extendable, modifiable, and upgradable. This flexible architecture ensures that FISK can readily adapt to evolving investigation, accommodate new models or methodologies, and integrate advancements in future analyses (for example, including the modelling of the galaxy bias).
@@ 4.1 Understanding the impact of the luminosity function parameters on the redshift distribution
We start by examining how changes in the parameters of the luminosity function affect the redshift distribution using the method described in Section 2.2. Understanding these effects will help us interpret later results and create realistic LSST-like redshift distributions as outlined in Section 3.
In our analysis, we employ the LF-based sample-PZ modelling methodology within the LSST survey framework. Our objective was to understand the influences of individual Schechter luminosity function parameters on the redshift distribution, a critical aspect for understanding results of our Fisher forecast further on. Each parameter was varied in isolation, with others held constant at their central values, enabling a focused investigation into the unique effects of each factor. The findings are shown in Fig. 5.
@@
    The normalization factor, $\phi_{0}^{*}$, which denotes the number density of galaxies at a benchmark luminosity and redshift, predominantly modulates the overall amplitude of the redshift distribution. This parameter acts as a scaling factor, and its variation is indicative of changes in the overall galaxy population density over cosmic time. The parametric sweep over the normalization parameter is indicated in green solid lines in Fig. 5. An increase in $\phi_{0}^{*}$ suggests a denser universe with a higher overall number of galaxies, amplifying the amplitude of the distribution.
@@
    In the context of the luminosity function’s influence on the redshift distribution of galaxies, the parameter $P$, which encapsulates the concept of galaxy density evolution with redshift, exerts a multifaceted impact. Specifically, $P$ modulates both the amplitude and the peak positioning of the redshift distribution. An increase in the $P$ value indicates a heightened density of galaxies at higher redshifts, as illustrated by the red solid lines in Figure 5. This rise in density not only boosts the total count of galaxies within the observed distribution but also leads to a notable shift in the distribution’s peak towards higher redshifts. Such a shift, facilitated by a higher $P$ value, describes a more compact arrangement of galaxies, leading to lower lacunarity within our fixed volume element.
@@
    In the analysis exploring the interplay between the luminosity function and the redshift distribution of galaxies, a key observation centers around the "knee" of the luminosity function, characterized by the parameter $M_{0}^{*}$. We have systematically increased $M_{0}^{*}$ from -22 to -18 in increments of 1, as shown in blue solid lines in Fig. 5. Notably, as the characteristic magnitude increased (indicating a shift towards dimmer characteristic luminosities), there is a marked decrease in the amplitude of the galaxy redshift distribution. Concurrently, a shift of the peak of the redshift distribution towards lower redshifts with larger $M_{0}^{*}$ values is present. This suggests that galaxies around and below the "knee" (dimmer galaxies) are more prevalent or detectable at lower redshifts, significantly influencing the shape and peak of the redshift distribution.
@@
    When modelling the redshift distribution of galaxies based on the luminosity function, an increase in the (fiducially negative) $\alpha$ value (indicating a flatter faint-end slope and hence a smaller number of low-luminosity galaxies) is observed to cause a decrease in the amplitude of the redshift distribution (distributions in purple in Fig. 5). Lower-luminosity galaxies make up a substantial portion of the overall galaxy number density across all redshifts. Although there is some corresponding increase in galaxies at luminosities just above the pivot, this effect is largely washed out by the dominance of the exponential term at higher luminosities. Since we have fixed the normalisation parameter $\phi_{0}^{*}$ for this exercise, it then follows that reducing the number of intrinsically faint galaxies will reduce the overall amplitude of the redshift distribution. Additionally, there is a noted shift in the peak of the redshift distribution towards higher redshifts as $\alpha$ increases towards zero. This shift could be attributed to a minor but still present increased relative representation of higher-luminosity galaxies (still in the faint-end i.e. below pivot luminosity), recalling that at higher redshifts we are preferentially going to observe intrinsically brighter galaxies.
In summary, these parameters collectively characterize the galaxy population, each contributing to the shape of the redshift distribution of galaxies over cosmic time. Furthermore, it is evident from the mathematical expressions (Eq. 4) that there exists a degree of degeneracy among these parameters. Such degeneracies are critical in understanding the limitations and constraints of our model and interpretations. The exact nature and directions of these degeneracies will be explored in the forthcoming discussion on our Fisher analysis results.
@@
Figure 6: Impact of Luminosity Function Parameters on Cosmic Shear Angular Power Spectra. This figure illustrates the relative difference in cosmic shear angular power spectra $C_{\ell}$, as a function of various Schechter luminosity function parameters. LF parameters are varied over five values, with the central value serving as the reference for relative comparison. In each plot, only the first bin correlations are displayed (bin 1- bin 1 correlation). The normalization factor $\phi_{0}^{*}$ is shown in green solid lines, highlighting that an increase leads to a negative relative difference ratio. The density evolution parameter $P$ is depicted in red solid lines, where an increase results in a higher density of galaxies at higher redshifts, thus augmenting the shear signal. The characteristic magnitude $M_{0}^{*}$ is in blue, and the rate of change $Q$ in orange-yellow, both showing a similar positive trend upon increase. Conversely, an increase in the faint-end slope $\alpha$ (in purple) is observed to also contribute to the shear signal, evident from the deviation pattern across different scales. The adjustment of each luminosity function parameter, though executed in equal incremental steps, does not uniformly translate to proportional changes in the amplitude of $C_{\ell}$. In some instances it is clear that as we deviate further from the central value, either by increasing or decreasing a parameter, the impact on the amplitude changes disproportionately. This non-linear response highlights the complex, non-proportional relationship between the variations in luminosity function parameters and their consequent effects on the cosmic shear strength.
@@
    Increasing the normalisation factor $\phi_{0}^{*}$, resulted in the relative difference ratio consistently being below zero (left top panel in Fig. 6). Physically, a higher $\phi_{0}^{*}$ implies an increased number density of galaxies with respect to luminosity. We attribute this effect to the impact of the LF parameter on the IA’s contribution to the spectrum. As a normalisation factor, the effect of $\phi_{0}^{*}$ will ultimately be normalised out of the lensing contribution. However, due to the way it is incorporated into the IA amplitude in our model (specifically, the fact that the red fraction, Eq. 11 contains a factor of $\phi_{0}^{*}$ for the red galaxy selection), this normalisation does not occur. So, an increase in $\phi_{0}^{*}$ increases the fraction of red galaxies and hence the IA contribution, and thus causes a relative decrease in the overall spectrum.
@@
    An increase in the density evolution parameter $P$ leads to an augmented relative difference ratio, as shown in the top middle panel in Fig. 6. This result can be interpreted as being primarily due to the lensing contribution to the spectrum, as increasing $P$ shifts the mean redshift of the tomographic bin up and therefore increases the amplitude of the lensing signal.
@@
    An increase in (fiducially negative) $\alpha$, the faint-end slope, also increases the shear signal, as can be seen in the bottom right panel in Fig. 6. This indicates that a shallower faint-end slope, corresponding to a relatively higher number of high-luminosity galaxies, ultimately contributes to the overall cosmic shear signal (and increases its amplitude).